An introduction to the Dutch L^{A}TEX document classes^{∗}

Victor Eijkhout

3 September 1989

1 Background

The Dutch L^{A}TEX document classes are an attempt to provide L^{A}TEX users with facilities of the standard distribution document classes ‘article’ and ‘report’, but coupled to typography that is geared towards Dutch usage. In order to ensure interchangeability, the new styles implemented the commands of the distribution styles. In principle, then, no manual for these styles would be needed. There are a few points, however, that can’t go without comment.

Table 1 lists all files connected with the Dutch document classes. These files can be found on the tex-nl fileserver on the hearn node of Bitnet/Earn and on the ctan. Not all of them are necessary, the ones with extension .tex contain just blah.

Remark: there exists a Dutch version of ‘letter’, which is so specifically Dutch that I decided not to include it in this list.

ntgclass.tex The L^{A}TEX 2ε source for this text
ntgclass.dtx The source for all the class files
ntgclass.dst A docstrip program to produce

the class files from the source.

artikel1.cls article compatible, design 1, straightforward
artikel2.cls article compatible, design 2, rather different
artikel3.cls article compatible, design 3, parskip instead of indent
rapport1.cls report compatible, design 1
rapport3.cls report compatible, design 3
book.cls book compatible, design 1
ntg10.clo 10 point option for all styles
ntg11.clo 11 point option for all styles
ntg12.clo 12 point option for all styles
artdoc.tex the genesis of the ‘artikel’ classes, in Dutch
briefdoc.tex the genesis of the ‘brief class, also in Dutch
rapdoc.tex the genesis of the ‘rapport’ classes, also in Dutch.

Table 1: List of files

∗ Updated for L^{A}TEX 2ε by Johannes Braams, 6 february 1994

2 Languages

These styles have been developed bearing in mind suggestions of Hubert Partl for making the language of styles switchable. Thus, on their own these styles will produce English captions like ‘chapter’ and ‘table of contents’, but specifying the options ‘dutch’ or ‘german’, or any language option that follows the directions of Partl, will switch these to the language of the option.

3 Design

At the moment there are styles compatible with ‘article’ and with ‘report’. The Dutch names for these are ‘artikel’ and ‘rapport’. In contrast to the standard styles, however, the user can now choose from different visual designs. Names of the styles are formed by appending their number to the name, for instance ‘rapport3’.

1. Design one is meant to be a universally acceptable style. It has been kept as uncontroversial as possible. Under heavy protests of the implementer (me) the one point that has turned out to be controversial, the table of contents, has been made subject to a switch that can restore the old L^{A}TEX layout. Explanation of this follows.

2. Design two will probably never be heavily used. It is more something of a heroic attempt to be different. At the moment only available in ‘artikel’ form.

3. Design three meets the wishes of people who like a zero \parindent and a positive \parskip. As just setting these parameters within for instance ‘artikel1’ will give some unwanted side effects, I decided to repair these in a separate style.

Credits for the visual design go to one real-life designer and a couple of books by designers I consulted. The full story can be found elsewhere.

3.1 User options As was mentioned above, the new layout of the table of contents has turned out to be somewhat controversial. So, in order to ensure wider acceptance of these styles I have incorporated a switch that will restore the old layout. Just specify the oldtoc option. This is available in designs 1 and 3. Number 2 really insists on being different.

3.2 Option files Like in the standard styles there exist option files for 10/11/12 point layout. Until now, however, I have managed to get away with using the same option files for both the ‘artikel’ ‘rapport’, and ‘boek’ styles. The option files then have to have some neutral names. Which are at the moment ntg10, ntg11, and ntg12. ‘NTG’, of course, stands for Nederlandse TEX Gebruikersgroep, i.e., Dutch TEX Users Group.

When doing the ‘rapport’ styles, I needed to modify the ‘titlepag’ option file. Thus there is a Dutch version of that, bearing the name of ‘voorwerk’ (literal translation: ‘frontwork’). Probably the majority of Dutch L^{A}TEX users don’t even know that this is the correct term. You’re never too old to learn.

With the L^{A}TEX 2ε version of thes classes the file voorwerk.sty has disappeared. It has been turned into the internal option voorwerk, just like titlpage.sty has disappeared into the internal option titlepage.